Curriculum Vitae et Studiorum

of

Francesco Piazza

• Born: .......................... 3^{rd} November 1970

• Nazionality: ................ italian

• Civil status: ................ celibate

• Military service: ......... free

• Permanent address: ... Via G.B. Sermei 9, 50014 Fiesole (FI) Tel. +39 (0)55 599624 – mob. (GSM) +39 349 8522727 • Office: ........................ IPT–BSP, EPFL, CH-1015 Lausanne, Switzerland. Tel. +41-(0)21-6930513

• E mail: ........................ piazza@fi.infn.it

• Webpage: .................... http://wwwpeople.unil.ch/Francesco.Piazza/

Present position

• From 1^{st} January 2003. Assistant Post–Doctoral in the Theoretical Biophysics Laboratory, Faculty of Sciences, École Polytechnique Fédérale de Lausanne (EPFL), Lausanne, Switzerland.

Past work experiences

• February 2001 – February 2003: INFM post–doc fellowship, FORUM project Study of statistical and dynamical models of biomolecules. Supervisor: prof. Roberto Livi, Dipartimento di Fisica, Università di Firenze.

Education

• October 1997 – October 2000. Heriot–Watt University, Edinburgh (UK). Ph.D. degree in Physics.

• 1989 – 1996: Università di Firenze. Physics Degree, final mark: 106/110. Laurea thesis’ title: High–T_{c} superconductors: study of electronic coupling through anharmonic vibrational modes.

• 1984 – 1989: Classical Lyceum “Dante Alighieri”, Florence. Diploma, mark: 60/60.

Teaching experiences

• September 2002 – February 2003: Lecturer at the University of Florence, Engineering faculty. General Physics I, first semester.

• July 2002: Invited lecture on “Statistical analysis of DNA sequences: a review”, International IP–Socrates summer school 2002, Villa Agape, Florence. Lecture available on–line at:

http://www.complex.unifi.it/WebComplex/bin/view/Time series analysis/TeachingMaterials

• October 1997 – October 1999: Heriot-Watt University, Physics Department. Tutorships for the following courses: Experimental Physics I e II (term 1–3 1997) and General Physics I (term 1–3 1998).

• 1994 – 1997: Mathematics and Science teacher at the intermediate school G. Barberi, Borgo Pinti, Florence.

• 1995 – 1997: Language support teacher for foreign children at the intermediate school G. Verdi, Via Monteverdi, Florence.

Other work experiences

• November 1999 – March 2001: Soprintendenza per i Beni Artistici e Storici di Firenze, Galleria Palatina di Palazzo Pitti, Firenze. Assistent technical manager (part–time).

• May – December 1996: Coordinator of the on–the–road social workers at the young–people centre Rione S. Croce (now young–people centre JAVA), Florence.

• July 1995: Animator at the local children summer centre, Impruneta (Florence).

• 1989 – 2001: Private tuitions to high–school and university students for the following subjects: Latin, Ancient greek, Mathematics and Physics.

Grants

• October 1997 – October 2000: Three–year scholarship from Heriot–Watt University, Edinburgh, towards the coverage of both living expenses and Ph.D. tuition fees.

Informatic knowledge

• Programming languages: FORTRAN, BASIC.

• Operating systems: Dos, Windows 98, Macintosh 9.0, VMS, UNIX (Linux).

• Word processors: Microsoft Word, L^{A}TEX.

• Spreadsheets: MS Excel, Jandel Sigma Plot, Kaleidagraph, Origin.

• Electronic editing systems: Aldus PageMaker, MS PowerPoint.

• Mathematics programs: Matlab, Maple V, Mathematica.

• Good knowledge of the programs: Adobe Illustrator, IDL, XMGR, GnuPlot.

Languages

• Excellent knowledge of written and spoken English.

• Elementary knowledge of written and spoken French.

Schools and Workshop

• 17 – 27 October 1999. Stage at ISI, Villa Gualino, Torino. Workshop “Complexity and Chaos” 1999.

• 29 August – 10 September 1999: 1999 EPSRC Theory of condensed matter Summer school, Coleg Harlech, North Wales. School’s subject: “High–T_{c} superconductivity”.

• 5 – 9 January 1998: 1998 Winter School on Superconductivity, Interdisciplinary Research Centre on Superconductivity (IRC), Cambridge. School’s subject: “Technological applications of high–T_{c} superconductivity.”.

• April 1996: crash course on “Communication skills”, organised by Punto Giovani di Firenze for the social workers of Rione S. Croce young–people centre.

Talks and posters at international conferences

• 17 – 21 June 2002, San Lorenzo de El Escorial, Madrid, talk at International Workshop Localisation and Energy transfer in Nonlinear Systems. Talk’s title “Cooling non–linear lattices till energy localisation”

• 23 – 27 March 2002, Arcachon (FR), poster at International workshop DNA in chromatin, at the frontiers of physics and biology. Poster’s title “Study of low–complexity DNA sequences in Eukaryotes and Prokaryotes”

• 13 – 17 June 2001. Talk at Eu–Us International Workshop on Discrete Breathers/Intrinsic Localized Vibrations, FORTH, Heraklion, Crete. Talk’s title “Slow energy relaxation and Localisation in 1D and 2D lattices”

Scientific publications

1. L. Cianchi, P. Moretti and F. Piazza, Superconducting cuprates: a simple model of coupling between electronic holes through apical ions, Phys. Lett. A. 246, 451 (1998).

2. F. Piazza, S. Lepri and S. Livi, Slow energy relaxation and localization in 1D lattices, J. Phys. A. 34, 9803 (2001).

3. F. Piazza et al., Study of atomic motions in EuBa_{2}Cu_{3}O_{7−δ} by Mössbauer and EXAFS spectroscopies., J. Superc.. 14, 675 (2001).

4. F. Piazza, Nonlinear Lattice Dynamics of High–T_{c} Superconductors, Ph.D. thesis, Heriot–Watt University, Edinburgh (2001).

5. F. Piazza, S. Lepri and R. Livi, Cooling nonlinear lattices toward localisation, invited paper, Chaos – Focus Issue Nonlinear localised modes; Fundamental Concepts and Applications, 13 (2), (2003).

6. F. Piazza, S. Lepri and R. Livi, Thermal activation of breathers in 2D nonlinear lattices, in Localisation and Energy transfer in Nonlinear Systems, Ed. L. V/’azquez, M. P. Zorzano and R. MacKay, World Scientific, Singapore (2003).

7. F. Piazza, A quantum perturbative pair distribution for determining interatomic potentials from EXAFS, J. Phys. C, 14 (45), 11623 (2002).

Other interests

• Dogged music and piano student.

• Drawing and painting.